OpenStudy (anonymous):

Part 1: Solve the inequality -3x + 4 less than or greater to x + 8 and show all of your work. You can use <= to represent the less than or equal to symbol. Part 2: Use complete sentences to describe the graph of the solution from Part 1.

OpenStudy (anonymous):

can someone please solve this??? I need to go visit my uncle in the hospital but i need to finish this first!

OpenStudy (anonymous):

-3x+4<=x+8 -3x-x<=8-4 -4x<=4 -x<= 1 * (-1) x>= -1 it means X= -1 or X> -1